OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 14 when x = 8. A. 112 B. 2 C. seven-fourths D. 7

OpenStudy (anonymous):

x is the coordinate to the left and y is to the right. so you start with x first

OpenStudy (anonymous):

you need to find out what with what makes 14

OpenStudy (anonymous):

if you add 2 three times to x then you get 14 but I don't think that's the answer

OpenStudy (paki):

y=k/x...as inverse... so 14=k/8 ==> k= 112... @musa19965